# TSD-SR: One-Step Diffusion with Target Score Distillation for Real-World Image Super-Resolution

### v3

@@ 3.4 Distribution-Aware Sampling Module
In the VSD-based framework, we need to match the score functions predicted by the Teacher Model with those by the LoRA Model at different timesteps $t\sim\{0,1,\cdots,T\}$. However, for the Real-ISR problem, this matching ability at different $t$ is unbalanced, as illustrated in Fig. 5(a) . This phenomenon could be attributed to our leveraging of the low-frequency (LF) priors from LQ data ,but losing guidance from high-frequency (HF) details. The noisy samples $\hat{z}_{t}$ possesses low-frequency priors (from LQ data). These priors are easily learned by the LoRA Model, leading to more similar predictions (Fig. 5(b)) during LF restoration (Stage 1). However, In the Stage 2, the lack of HF details in $\hat{z}_{t}$ causes the LoRA Model to poorly perceive details, resulting in divergent predictions (Fig. 5(c)). Our goal is to reduce this divergence.
@@
Existing method, for each iteration, only matches the score function at a single timestep sample $\hat{z}_{t}$, with uniform sampling strategy of time $t$. This leads to slow convergence or even difficulty in optimization during Stage 2, as the gradients of important timesteps are averaged out. To this end, we propose our Distribution-Aware Sampling Module (DASM). This module accumulates optimization gradients for earlier timestep samples in a single iteration, enabling the backpropagation of more gradients focused on detail optimization. As shown in Fig. 6, we first introduce noise into the synthetic latent representation $\boldsymbol{\hat{z}_{t}}=(1-\sigma_{t})\boldsymbol{\hat{z}_{0}}+\sigma_{t}\boldsymbol{\epsilon}$, where $\sigma$ is a weight factor and $\epsilon$ is a gaussian noise. Subsequently, we employ a LoRA Model to perform denoising, yielding the previous time noisy samples by Eq. 6:
@@
where $\sigma_{t-1}$ and $\sigma_{t}$ are from the flow matching scheduler, and here LoRA Model has fit $\boldsymbol{\hat{z}_{0}}$’s distribution. Similarly, $\boldsymbol{z_{t-1}}$ can be obtained by denoising on the Teacher Model. Ultimately, in a single iteration, the gradients from noise samples along its trajectory can be accumulated to update the Student Model. Since the obtained samples follow the diffusion sampling trajectory and are directed towards early timestep, we effectively optimize the divergence in Stage 2.
Figure 6: Illustration of DASM. Top: The naive method that add noise to samples. Bottom: Our DASM leverages the priors of diffusion models to sample noise that more closely aligns with the true sampling trajectory. The noisy samples can all serve as inputs to the subsequent network to generate gradient backpropagation.
@@ 3.5 Training Objective
For the reconstruction loss, we use $LPIPS$ loss and $MSE$ loss in the latent space:
$\begin{split}&\mathcal{L}_{Rec}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right)\\ &=\gamma_{1}\mathcal{L}_{LPIPS}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right)+\mathcal{L}_{MSE}\left(\mathbf{z}_{t},\mathbf{\hat{z}}_{t}\right).\end{split}$ (7)
For the regularization loss, we use our TSD loss Eq. 5. Therefore, the overall training objective for the Student Model $G_{\theta}$ is:
@@
LoRA Model $\boldsymbol{\epsilon}_{\phi}$. As stipulated by VSD, the fine-tuned replica $\epsilon_{\phi}$ must be trainable, with its training objective being:
@@
where $\boldsymbol{\epsilon}^{\prime}$ serves as the training target for the denoising network, representing Gaussian noise in the context of DDPM, and a gradient towards HQ for flow matching.
@@ 4.1 Experimental Settings
Training Datasets. We adopt DIV2K [1], Flickr2K [43], LSDIR [27] and first 10K face images from FFHQ [20] for training. We use the same degradation pipeline as Real-ESRGAN [47] to synthesize LR-HR pairs.
Test Datasets. We evaluated our model on synthetic DIV2K-Val [1] dataset and two real-world datasets, RealSR [4] and DRealSR [52]. The real-world datasets comprised 128x128 LQ and 512x512 HQ image pairs. The synthetic dataset included 3,000 image pairs, which are created by downgrading the 512x512 HQ images croped from DIV2K-Val to 128x128 LQ images using the Real-ESRGAN [47] degradation pipeline.
Evaluation Metrics. For evaluating our method, we apply both full-reference and no-reference metrics. Full-reference metrics include PSNR and SSIM [50] (calculated on the Y channel in YCbCr space) for fidelity, LPIPS [66] and DISTS [7] for perceptual quality, and FID [14] for distribution comparison. No-reference metrics include NIQE [64], MANIQA [56], MUSIQ [22] and CLIPIQA [44].
@@
Implementation Details. All models mentioned are initialized from the Teacher Model (SD3 [10] in our work). Similar to OSEDiff [53], we only train the vae encoder and denoising network in the Student Model, frozen the vae decoder in order to preserve vae’s prior [24]. We utilize the default prompt for the Student Model, and prompts are extracted from HQ for the Teacher and LoRA models when training. We use AdamW optimizer [30] with a learning rate of 5e-6 for the Student Model and 1e-6 for the LoRA Model, and set the rank of LoRA to 64 for both two models. During the initial training phase, we incorporate MSE loss in the latent space and exclude DASM to enhance model stability and reduce training time. In later stages, we remove this MSE loss as it leads to over-smoothed results, and we employ DASM to improve image restoration. The training process took roughly 96h, utilizing 8 NVIDIA V100 GPUs with a batch size of 16.
@@ 4.2 Comparison with Existing Methods
Quantitative Comparisons. Tab. 1 show the quantitative comparison of our method with single-step diffusion models on three datasets. Ours achieves the best results on most evaluation metrics. SinSR and AddSR, as distilled versions of previous multi-step super-resolution methods, reduce inference steps but experience a corresponding decrease in performance metrics. OSEDiff introduces the VSD loss from 3D generation tasks into Real-ISR without fully addressing the substantial differences between the two domains. Therefore, the no-reference metrics for image quality are not very satisfied. In contrast, our proposed TSD-SR, tailored for Real-ISR, outperforms other single-step models in terms of the vast majority of key metrics.
Tab. 2 show the quantitative comparison with multi-step models. We can draw the following conclusions: (1) TSD-SR demonstrates significant advantages over competing methods in terms of LPIPS, DISTS, and NIQE metrics. Additionally, it achieves performance that surpasses most multi-step models on FID, MUSIQ, and CLIPIQA. (2) DiffBIR, SeeSR, PASD, and AddSR exhibit better performance on the MANIQA metric, which may be attributed to the fact that multi-step models have more denoising iterations to produce rich details. (3) ResShift stands out with the highest PSNR and SSIM scores, while StableSR also shows notable performance regarding DISTS and FID metrics. However, both two models underperform in terms of no-reference metrics.
Finally, we explain the lower PSNR and SSIM metrics in our experiments. Several works [55, 61] have found that these reconstruction metrics are not well-suited for evaluating Real-ISR tasks. When a model recovers better details, it leads to lower reconstruction metrics, indicating a trade-off [3, 68, 32]. This phenomenon has also been widely discussed in other related research work [3, 68, 32, 61, 66, 45, 55]. LPIPS [66] is proposed to overcome the limitation that PSNR and SSIM fail to align with human judgments in spatial ambiguities situation. Other DMs-based SR researchers [61, 45] argue that DMs introduce superior pre-trained priors, enabling the restoration of information that traditional methods (from scratch) cannot achieve. However, such capability often leads to a decline in pixel-level metrics, as they prioritize distribution modeling and sampling from learned distributions over strict pixel fidelity. We anticipate the development of better full-reference metrics in the future to assess advanced Real-ISR methods. Refer to the supplementary material for detailed visual comparisons.
@@
Qualitative Comparisons. Fig. 7 present visual comparisons of different Real-ISR methods. As shown in multi-step methods’ results, SeeSR uses degradation-aware semantic cues to leverage image generation priors, but it sometimes produces over-smooth textures. SUPIR demonstrates a notably robust generative capacity, however, the overproduction of extraneous details will lead to results that lack naturalism in image restoration (e.g. Adding excessive wrinkles at the corners of a young girl’s eyes). Under more realistic degradation conditions, PASD finds it challenging to restore the corresponding content, revealing a deficiency in the model’s robustness capabilities. Among single-step methods, SinSR introduces artifacts, possibly because its distilled pre-trained model is trained from scratch, lacking sufficient real-world priors, which results in unsatisfactory image restoration capabilities. AddSR produces over-smoothed results when using its 1-step model. OSEDiff offers improved restoration effects compared to SinSR and AddSR, yet it may fall short in terms of authenticity and naturalness when it comes to detail recovery. In contrast, our method effectively generates rich textures and realistic details with enhanced sharpness and contrast. Additional visual comparisons and results are provided in the supplementary material.
@@
Table 3: Comparison of computational complexity across different DMs-based methods. The performance of each method is measured on an A100 GPU, using input images sized at 512 × 512 pixels. We have disregarded the loading time for model weights and data.
@@
We assess the complexity of the state-of-the-art DM-based Real-ISR models as detailed in Tab. 3, focusing on inference time. The performance of each method is benchmarked on an A100 GPU with input images sized at 512 × 512 pixels. We have disregarded the loading time for model weights and data, and the main computation time includes: (1) text extraction time when a text extractor is used; (2) text encoder computation time if needed; (3) VAE encoding and decoding time; (4) denoising network execution time. It is evident that TSD-SR has a substantial advantage in inference time compared with multi-step models. Specifically, TSD-SR is more than 120 times faster than SUPIR, 90 times faster than StableSR, about 50 times faster than DiffBIR, more than 40 times faster than SeeSR, more than 35 times faster than PASD, and 4 times faster than ResShift. When compared with existing one-step models, our model boasts the fastest inference times. This is because we directly denoise from LQ data and use a fixed prompt.
@@ 4.3 User Study
We conduct a user study comparing our method with three other diffusion-based one-step super-resolution methods. To make the evaluation more comprehensive, we selected images from five categories—human faces, buildings, animals, vegetation and characters. A total of 50 participants were engaged in the voting process. We guided participants to evaluate the best restoration results based on the similarity to HQ image, structural similarity to the LQ image, and realism of textures and details. The results shown in the Fig. 8 indicate that our method receives a 69.2% approval rate from users. We scored 57.60% in Animals, 70.00% in Buildings, 68.80% in Human Faces, 65.20% in Vegetation, and 84.40% in Characters, surpassing other models.
Figure 8: The results of our user study. Left: Category-based user preference radar chart. Our model has won the most user favor across all categories. Right: User preference pie chart. Our approach has garnered a 69.2% user satisfaction rating.
@@ 4.4 Ablation Study
Effectiveness of TSM and DASM. To validate the effectiveness of TSM loss and DASM, we conduct ablation studies by removing them separately in our experiments. We select LPIPS, DISTS, MUSIQ, MANIQA and CLIPIQA for comparison, as these metrics are critical for image quality assessment. Additionally, we select FID as the metric to evaluate distribution similarity. The results are presented in Tab. 4. We draw the following conclusions: (1) The absence of TSM loss and DASM negatively impacts performance across both reference-based (LPIPS, DISTS) and non-reference-based metrics (MUSIQ, MANIQA and CLIPIQA). The FID metric is also adversely affected, suggesting a decline in distribution fidelity. (2) The lack of TSM leads to a significant decrease in LPIPS, DISTS, MUSIQ, and CLIPIQA metrics, possibly due to the unreliable directions in VSD causing unrealistic generations. (3) The absence of DASM results in a decline in FID, MUSIQ and CLIPIQA, possibly due to the suboptimal optimization of details.
@@
Base model for fairer comparison. To validate the effectiveness of our method across different versions of SD models, we conduct additional experiments, as shown in Tab. 5, including SD2-base and SD2.1-base models. The performance is evaluated on the DRealSR test dataset [52]. It demonstrates superior performance compared to other one-step SR methods, including OSEDiff [53] and AddSR [55]. Specifically, our SD2-base version model outperforms the single-step AddSR across all perceptual reference metrics and no-reference metrics, particularly excelling in NIQE[64], MUSIQ[22], and CLIPIQA[44], significantly surpassing the performance of one step AddSR. Our SD2.1-base model demonstrates comparable performance to OSEDiff and surpasses it across various metrics, with notable improvements in NIQE and CLIPIQA.
Table 5: Same base model for fairer comparison to validate the effectiveness of TSD-SR.
@@
Parameters $N$ and $s$ in DASM. We provide performance comparisons for different combinations of $N$ and $s$ in Tab. 6. The performance is evaluated on the DRealSR test dataset. $N$ is set to 4, and $s$ is set to 50 in our setting (bold in the table). Larger or smaller $N$ will degrade performance, possibly because it is related to regularization strength. We prefer to use a smaller $N$ because DASM is computationally time-consuming. Therefore, after carefully balancing training duration and model performance, we select $N$=4 as the final value. Small $s$ will have similar performance, but the image quality will be compromised when setting large $s$. Our experimental results indicate that selecting $s$ within the range of 25-75 may yield better performance.

### v4

@@ 3.4 Distribution-Aware Sampling Module
In the VSD-based framework, it is necessary to match the score functions predicted by the Teacher Model and the LoRA Model across timesteps $t\in{0,1,\dots,T}$. However, for the Real-ISR problem, this matching performance is inconsistent across timesteps, as illustrated in Fig. 5(a). This phenomenon may be attributed to the reliance on low-frequency (LF) priors in the LQ data, while lacking guidance from high-frequency (HF) details. The output sample $\boldsymbol{\hat{z}_{0}}$, derived from LQ data, contains low-frequency (LF) priors that are easily captured by the LoRA Model, resulting in similar predictions during LF restoration (Stage 1), as shown in Fig. 5(b). However, in Stage 2, due to the absence of high-frequency (HF) details in $\boldsymbol{\hat{z}_{0}}$, the LoRA Model struggles to reconstruct fine-grained features, leading to divergent predictions, as illustrated in Fig. 5(c). To address this issue, we aim to reduce such divergence.
@@
Existing methods match the score function at each iteration using a single latent sample $\boldsymbol{\hat{z}_{t}}$, with the timestep $t$ drawn from a uniform distribution. This leads to slow convergence and even training difficulty during Stage 2, as gradients from important timesteps are diluted by uniform averaging. To this end, we propose our Distribution-Aware Sampling Module (DASM). This module accumulates optimization gradients for earlier timestep samples in a single iteration, enabling the backpropagation of more gradients focused on detail optimization. As shown in Fig. 6, we first obtain the noisy synthetic latent representation as $\boldsymbol{\hat{z}_{t}}=(1-\sigma_{t})\boldsymbol{\hat{z}_{0}}+\sigma_{t}\boldsymbol{\epsilon}$, where $\sigma_{t}$ is a weighting factor and $\epsilon$ denotes Gaussian noise. Subsequently, we employ a LoRA Model to perform denoising, yielding noisy samples at the previous timestep as described in Eq. 6:
@@
The parameters $\sigma_{t-1}$ and $\sigma_{t}$ are obtained from the flow matching scheduler. Here, the LoRA Model has learned the distribution of $\boldsymbol{\hat{z}_{0}}$. Similarly, $\boldsymbol{z_{t-1}}$ can be obtained by denoising using the Teacher Model. In a single iteration, gradients from noisy samples along the sampling trajectory can be accumulated to update the Student Model. Since these samples follow the diffusion sampling trajectory and are concentrated at early timesteps, this approach effectively reduces the divergence observed in Stage 2.
Figure 6: Illustration of DASM. Top: The naive approach that adds noise directly to the samples. Bottom: The proposed DASM leverages diffusion model priors to generate noisy latent that better align with the true sampling trajectory. These noisy samples can all serve as inputs to the downstream network, enabling effective gradient backpropagation.
@@ 3.5 Training Objective
For the reconstruction loss, we use the $LPIPS$ loss in the pixel space and the $MSE$ loss in the latent space:
$\begin{split}&\mathcal{L}_{Rec}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right)\\ &=\gamma_{1}\mathcal{L}_{LPIPS}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right)+\mathcal{L}_{MSE}\left(\boldsymbol{z}_{t},\boldsymbol{\hat{z}}_{t}\right).\end{split}$ (7)
For the regularization loss, we use our TSD loss, Eq. 5. Therefore, the overall training objective for the Student Model $G_{\theta}$ is:
@@
LoRA Model $\boldsymbol{\epsilon}_{\phi}$. As stipulated by VSD, the replica $\epsilon_{\phi}$ must be trainable, with its training objective being:
@@
where $\boldsymbol{\epsilon}^{\prime}$ serves as the training target for the denoising network, representing Gaussian noise in the context of DDPM, and a gradient towards HQ data for flow matching.
@@ 4.1 Experimental Settings
Training Datasets. For training, we utilize DIV2K [1], Flickr2K [43], LSDIR [27], and the first 10K face images from FFHQ [20]. To synthesize LR-HR pairs, we adopt the same degradation pipeline as in Real-ESRGAN [47].
Test Datasets. We evaluate our model on the synthetic DIV2K-Val [1] dataset, as well as two real-world datasets: RealSR [4] and DRealSR [52]. The real-world datasets consist of 128×128 low-quality (LQ) and 512×512 high-quality (HQ) image pairs. For the synthetic set, 3,000 pairs were generated by cropping 512×512 patches from DIV2K-Val and applying the Real-ESRGAN [47] degradation pipeline to downsample them to 128×128.
Evaluation Metrics. To evaluate our method, we employ both full-reference and no-reference metrics. The full-reference metrics include PSNR and SSIM [50] (computed on the Y channel of the YCbCr color space) for fidelity; LPIPS [66] and DISTS [7] for perceptual quality; and FID [14] for measuring distribution similarity. The no-reference metrics include NIQE [64], MANIQA [56], MUSIQ [22], and CLIPIQA [44].
@@
Implementation Details. All models are initialized from the Teacher Model (SD3 [10] in our work). Similar to OSEDiff [53], we only train the VAE encoder and the denoising network in the Student Model, freezing the VAE decoder to preserve its prior [24]. We utilize the default prompt for the Student Model, while prompts are extracted from HQ images for the Teacher and LoRA models during training. We adopt the AdamW optimizer [30] with a learning rate of 5e-6 for the Student Model and 1e-6 for the LoRA Model, setting the LoRA rank to 64 for both models. During the initial training phase, we incorporate MSE loss in the latent space and exclude DASM to stabilize training and reduce time cost. In later stages, we remove the MSE loss to avoid over-smoothed results and introduce DASM to enhance restoration quality. The training process took approximately 96 hours, utilizing 8 NVIDIA V100 GPUs with a batch size of 16.
@@ 4.2 Comparison with Existing Methods
Quantitative Comparisons. Tab. 1 shows the quantitative comparison of our method with other single-step diffusion models on three datasets. Our method achieves the best results across most evaluation metrics. SinSR and AddSR, as distilled versions of previous multi-step super-resolution methods, significantly reduce inference steps while suffering a corresponding drop in performance. OSEDiff introduces the VSD loss from 3D generation tasks into the Real-ISR without fully accounting for the substantial differences between the two domains. As a result, its no-reference image quality metrics are not satisfactory. In contrast, our proposed TSD-SR, specifically designed for Real-ISR, outperforms all other single-step models in the vast majority of key metrics.
Tab. 2 shows the quantitative comparison with multi-step models. We can draw the following conclusions: (1) TSD-SR demonstrates significant advantages over competing methods in terms of LPIPS, DISTS, and NIQE. Additionally, it outperforms most multi-step models in FID, MUSIQ, and CLIPIQA. (2) DiffBIR, SeeSR, PASD, and AddSR achieve better results in terms of MANIQA, which may be attributed to the fact that multi-step models benefit from more denoising iterations to generate richer details. (3) ResShift stands out with the highest PSNR and SSIM scores, while StableSR also performs well in terms of DISTS and FID. However, both models underperform on the no-reference metrics.
Finally, we explain the relatively lower PSNR and SSIM scores observed in our experiments. Several studies [55, 61] have shown that these reconstruction metrics are not well-suited for the evaluation of Real-ISR tasks. Models that recover more realistic or detailed textures often yield lower PSNR and SSIM scores, reflecting a fundamental trade-off between perceptual quality and pixel-wise fidelity [3, 68, 32]. This phenomenon has also been extensively discussed in the other research work [3, 68, 32, 61, 66, 45, 55]. LPIPS [66] is proposed to overcome the limitation that PSNR and SSIM fail to align with human judgments in spatial ambiguities situations. Other DMs-based SR researchers [61, 45] argue that DMs introduce superior pre-trained priors, enabling the restoration of information that traditional methods (from scratch) cannot achieve. However, such capability often leads to a decline in pixel-level metrics, as they prioritize distribution modeling and sampling from learned distributions over strict pixel fidelity. We anticipate the development of better full-reference metrics in the future to assess advanced Real-ISR methods. Refer to the supplementary material for detailed visual comparisons.
@@
Qualitative Comparisons. Fig. 7 presents visual comparisons of different Real-ISR methods. As shown in the results of multi-step methods, SeeSR leverages degradation-aware semantic cues to incorporate image generation priors, but it tends to produce over-smoothed textures in some cases. SUPIR demonstrates notably robust generative capabilities. However, the excessive generation of fine details can result in outputs that appear less natural (e.g., adding unnecessary wrinkles around the eyes of a young girl). Under more realistic degradation conditions, PASD finds it difficult to recover the appropriate content, indicating limited robustness. Among single-step methods, SinSR tends to produce artifacts, likely due to its base model, ResShift, being trained from scratch without adequate exposure to real-world priors, which leads to inferior image restoration quality. AddSR produces over-smoothed results when using its 1-step model. OSEDiff demonstrates better restoration performance than SinSR and AddSR; however, it may fall short in terms of authenticity and naturalness, particularly in recovering fine details. In contrast, our method effectively generates rich textures and realistic details with enhanced sharpness and contrast. Additional visual comparisons and results are provided in the supplementary material.
@@
Table 3: Comparison of computational complexity across different diffusion model-based methods. Performance is measured on an A100 GPU using 512×512 input images, excluding model weight and data loading time.
@@
We assess the computational complexity of the state-of-the-art DM-based Real-ISR methods, as detailed in Tab. 3, with a focus on inference time. Each method is benchmarked on an A100 GPU using input images of size 512×512 pixels. We disregarded the loading time for model weights and data. The main computation time consists of: (1) text extraction time (if a text extractor is used); (2) text encoder computation time (if applicable); (3) VAE encoding and decoding time; and (4) denoising network execution time. It is evident that TSD-SR holds a substantial advantage in inference speed compared to multi-step models. Specifically, TSD-SR is over 120× faster than SUPIR, 90× faster than StableSR, approximately 50× faster than DiffBIR, over 40× faster than SeeSR, more than 35× faster than PASD, and 4× faster than ResShift. When compared with existing one-step models, our method achieves the fastest inference times. This advantage is attributed to directly denoising from LQ data and employing a fixed prompt.
@@ 4.3 User Study
We conduct a user study comparing our method with three other diffusion-based one-step super-resolution methods. To ensure a comprehensive evaluation, we selected images from five categories—human faces, buildings, animals, vegetation, and characters. A total of 50 participants took part in the voting process. Participants were instructed to select the best restoration results based on similarity to the HQ image, structural similarity to the LQ image, and realism of textures and details. The results in Fig. 8 indicate that our method received a 69.2% approval rate from users. Specifically, our method achieved 57.6% in Animals, 70.0% in Buildings, 68.8% in Human Faces, 65.2% in Vegetation, and 84.4% in Characters, surpassing those of other methods.
Figure 8: Results of our user study. Left: Category-based user preference radar chart, showing that our model received the highest favor across all categories. Right: User preference pie chart, illustrating that our approach garnered a 69.2% user satisfaction rating.
@@ 4.4 Ablation Study
Effectiveness of TSM and DASM. To validate the effectiveness of the TSM loss and DASM, we conduct ablation studies by removing each component separately. We select LPIPS, DISTS, MUSIQ, MANIQA, and CLIPIQA for comparison, as these metrics are critical for image quality assessment. Additionally, FID is used to evaluate distribution similarity. The results are presented in Tab. 4. From the results, we draw the following conclusions: (1) The absence of TSM loss and DASM negatively impacts performance across both reference-based metrics (LPIPS, DISTS) and no-reference metrics (MUSIQ, MANIQA, and CLIPIQA). The FID metric is also adversely affected, indicating a decline in distribution fidelity. (2) The lack of TSM leads to a significant decrease in LPIPS, DISTS, MUSIQ, and CLIPIQA, possibly due to unreliable directions in VSD leading to unrealistic generations. (3) The absence of DASM results in a decline in FID, MUSIQ, and CLIPIQA, possibly due to suboptimal detail optimization.
@@
Base model for fairer comparison. To validate the effectiveness of our method across different versions of SD models, we conduct additional experiments on SD2-base and SD2.1-base models, as shown in Tab. 5. The performance is evaluated on the DRealSR test dataset [52]. Our method demonstrates superior performance compared to other one-step SR methods, including OSEDiff [53] and AddSR [55]. Specifically, our SD2-base model outperforms single-step AddSR across all perceptual reference and no-reference metrics, particularly excelling in NIQE [64], MUSIQ [22], and CLIPIQA [44]. Meanwhile, our SD2.1-base model shows comparable or better performance than OSEDiff across various metrics, with notable improvements in NIQE and CLIPIQA.
Table 5: Fair comparison using the same base model to validate TSD-SR
@@
Parameters $N$ and $s$ in DASM. We compare performance under different combinations of $N$ and $s$ in Tab. 6. The evaluation is conducted on the DRealSR test dataset. In our setting, $N$ is set to 4 and $s$ to 50 (highlighted in bold in the table). Performance degrades when $N$ is either larger or smaller, possibly due to its effect on regularization strength. Since DASM is computationally expensive, we prefer a smaller $N$. After balancing training time and performance, we select $N=4$ as the final value. Smaller values of $s$ yield similar performance, while larger values degrade image quality. Experimental results suggest that choosing $s$ between 25 and 75 achieves better overall performance.